# An Efficient GPU-based Implementation for Noise Robust Sound Source Localization

### v1

@@ Abstract
Robot audition, encompassing Sound Source Localization (SSL), Sound Source Separation (SSS), and Automatic Speech Recognition (ASR), enables robots and smart devices to acquire auditory capabilities similar to human hearing. Despite their wide applicability, processing multi-channel audio signals from microphone arrays in SSL involves computationally intensive matrix operations, which can hinder efficient deployment on Central Processing Units (CPUs), particularly in embedded systems with limited CPU resources. This paper introduces a GPU-based implementation of SSL for robot audition, utilizing the Generalized Singular Value Decomposition-based Multiple Signal Classification (GSVD-MUSIC), a noise-robust algorithm, within the HARK platform, an open-source software suite. For a 60-channel microphone array, the proposed implementation achieves significant performance improvements. On the Jetson AGX Orin, an embedded device powered by an NVIDIA GPU and ARM®Cortex-A78AE v8.2 64-bit CPUs, we observe speedups of 4645.1$\times$ for GSVD calculations and 8.8$\times$ for the SSL module, while speedups of 2223.4$\times$ for GSVD calculation and 8.95$\times$ for the entire SSL module on a server configured with an NVIDIA A100 GPU and AMD EPYC 7352 CPUs, making real-time processing feasible for large-scale microphone arrays and providing ample capacity for real-time processing of potential subsequent machine learning or deep learning tasks.
@@ I Introduction
To address these challenges, an efficient implementation is essential for real-time processing of large-scale microphone arrays. This paper presents a GPU-based implementation of SSL using Generalized Singular Value Decomposition-based Multiple Signal Classification (GSVD-MUSIC), a noise-robust algorithm, within HARK [13, 30], the leading open-source platform for robot audition. Our approach aims to enhance the applicability of robot audition in embedded scenarios with low signal-to-noise ratio (SNR) environments. For a 60-channel microphone array, on the Jetson AGX Orin—an embedded device powered by an NVIDIA GPU and ARM®Cortex-A78AE v8.2 64-bit CPUs—speedups of 4645.1$\times$ for the GSVD part and 8.8$\times$ for the SSL module are observed, while speedups of 2223.4$\times$ for GSVD calculation and 8.95$\times$ for the entire SSL module are obtained on a server with an NVIDIA A100 GPU and AMD EPYC 7352 CPUs. Our work enables real-time processing for large-scale microphone arrays and reserve spaces for achieving real-time subsequent ML/DL tasks.
@@ V-D Performance Results
Table I presents the performance evaluation results. The GPU-based implementation significantly outperforms the CPU-based implementation in terms of processing time. Specifically, on the Jetson AGX Orin, we observed speedups of 4645.1$\times$ for the GSVD calculation and 8.8$\times$ for the entire SSL module. On the A100 server, speedups of 2223.4$\times$ and 8.95$\times$ were observed for the GSVD calculation and the entire SSL module, respectively.
@@ VI Conclusion
This paper presents GPU implementations for GSVD-based SSL, a noise-robust algorithm, with a particular focus on deploying HARK, an open-source platform for robot audition. The proposed implementation significantly reduces processing times, enabling real-time operation with a 60-channel microphone array. Specifically, on the Jetson AGX Orin, speedups of 4645.1$\times$ and 8.8$\times$ were observed for the GSVD calculation and the entire SSL module, respectively, while speedups of 2223.4$\times$ for GSVD calculation and 8.95$\times$ for the entire SSL module are obtained on a server with an NVIDIA A100 GPU and AMD EPYC 7352 CPUs.

### v2

@@ Abstract
Robot audition, encompassing Sound Source Localization (SSL), Sound Source Separation (SSS), and Automatic Speech Recognition (ASR), enables robots and smart devices to acquire auditory capabilities similar to human hearing. Despite their wide applicability, processing multi-channel audio signals from microphone arrays in SSL involves computationally intensive matrix operations, which can hinder efficient deployment on Central Processing Units (CPUs), particularly in embedded systems with limited CPU resources. This paper introduces a GPU-based implementation of SSL for robot audition, utilizing the Generalized Singular Value Decomposition-based Multiple Signal Classification (GSVD-MUSIC), a noise-robust algorithm, within the HARK platform, an open-source software suite. For a 60-channel microphone array, the proposed implementation achieves significant performance improvements. On the Jetson AGX Orin, an embedded device powered by an NVIDIA GPU and ARM®Cortex-A78AE v8.2 64-bit CPUs, we observe speedups of 5648.7$\times$ for GSVD calculations and 10.7$\times$ for the SSL module, while speedups of 4245.1$\times$ for GSVD calculation and 17.3$\times$ for the entire SSL module on a server configured with an NVIDIA A100 GPU and AMD EPYC 7352 CPUs, making real-time processing feasible for large-scale microphone arrays and providing ample capacity for real-time processing of potential subsequent machine learning or deep leraning tasks.
@@ I Introduction
To address these challenges, an efficient implementation is essential for real-time processing of large-scale microphone arrays. This paper presents a GPU-based implementation of SSL using Generalized Singular Value Decomposition-based Multiple Signal Classification (GSVD-MUSIC), a noise-robust algorithm, within HARK [13, 30], the leading open-source platform for robot audition. Our approach aims to enhance the applicability of robot audition in embedded scenarios with low signal-to-noise ratio (SNR) environments. For a 60-channel microphone array, on the Jetson AGX Orin—an embedded device powered by an NVIDIA GPU and ARM®Cortex-A78AE v8.2 64-bit CPUs—speedups of 5648.7$\times$ for the GSVD part and 10.7$\times$ for the SSL module are observed, while speedups of 4245.1$\times$ for GSVD calculation and 17.3$\times$ for the entire SSL module are obtained on a server with an NVIDIA A100 GPU and AMD EPYC 7352 CPUs. Our work enables real-time processing for large-scale microphone arrays and reserve spaces for achieving real-time subsequent ML/DL tasks.
@@ V-D Performance Results
Table I presents the performance evaluation results. The GPU-based implementation significantly outperforms the CPU-based implementation in terms of processing time. Specifically, on the Jetson AGX Orin, we observed speedups of 5648.7$\times$ for the GSVD calculation and 10.7$\times$ for the entire SSL module. On the A100 server, speedups of 4245.1$\times$ and 17.3$\times$ were observed for the GSVD calculation and the entire SSL module, respectively.
@@ VI Conclusion
This paper presents GPU implementations for GSVD-based SSL, a noise-robust algorithm, with a particular focus on deploying HARK, an open-source platform for robot audition. The proposed implementation significantly reduces processing times, enabling real-time operation with a 60-channel microphone array. Specifically, on the Jetson AGX Orin, speedups of 5648.7$\times$ and 10.7$\times$ were observed for the GSVD calculation and the entire SSL module, respectively, while speedups of 4245.1$\times$ for GSVD calculation and 17.3$\times$ for the entire SSL module are obtained on a server with an NVIDIA A100 GPU and AMD EPYC 7352 CPUs.